OpenStudy (anonymous):

(5 radical 25d) (2d radical 3) What is the product? Simplify if possible.

OpenStudy (anonymous):

\[5*\sqrt{25d}*2d*\sqrt{3}\] or \[\sqrt[5]{25d} * \sqrt[2d]{3}\] ?

OpenStudy (anonymous):

whats the answer then

OpenStudy (anonymous):

If the first one, that is much easier, and I'm guessing that's the one you're asking about. If so, because the radicals are both square roots (not cube roots or anything more complex) we can multiply "like terms"... Let's simplify first. The first term is \[5\sqrt{25d}\] Since can take the sqrt of 25 (which is 5), we will do so and pull that number out front and multiply it by the coefficient that is already there to get \[5*5\sqrt{d}\] The second term cannot be simplified. Let's take the coefficients for both terms, (25 and 2d) and multiply them to get 50d (these are "like terms" because they are the coefficients out front of the radicals). Similarly, we can take what's under the radicals and combine them under one radical by multiplying,\[\sqrt{d}*\sqrt{3}= \sqrt{3d}\] Putting this all together, we get \[50d \sqrt{3d}\] HTH!

OpenStudy (anonymous):

k thx